# Asymptotics of the maximum likelihood estimator of the location parameter of Pearson Type VII distribution

### v1

@@ Abstract.
Pearson Type VII distribution, Cauchy distribution, maximum likelihood estimator, strong consistency, asymptotic efficiency, Bahadur efficiency
@@ 1. Introduction
The family of Pearson Type VII distributions provides flexible heavy-tailed models. The estimation of its parameters dates back at least to Fisher [7], nearly a century ago, and many researchers have studied it since then; see Johnson, Kotz, and Balakrishnan [9, Section 28] for a thorough survey of results prior to 1994. This class is also known as the location–scale family of Student’s $t$ distributions or of $q$-Gaussian distributions. For estimating the location, the median is a robust alternative to the arithmetic mean; however, it is not asymptotically efficient.
@@
In general, the maximum likelihood estimator is widely regarded as optimal in large samples under standard regularity. Lange, Little, and Taylor [10] proposed a strategy based on maximum likelihood for a general model with errors following the $t$-distribution and applied it to many problems. Under suitable regularity conditions, properties such as strong consistency, asymptotic efficiency, and Bahadur efficiency have been established by many researchers. For location–scale families, it is natural to consider the estimation of the location with known scale. The standard approach is to solve the likelihood equation explicitly or numerically, which often has a unique root. For the Cauchy distribution with known scale, however, the likelihood equation may have multiple roots (see Reeds [12] for precise analysis), and the same phenomenon occurs for the Pearson Type VII distribution. For this reason, alternative estimators of the Cauchy location parameter have been considered. For example, Cohen Freue [6] considered the Pitman estimator for small samples, and Zhang [17] considered an empirical Bayes estimator. Nevertheless, this does not represent a failure of the maximum likelihood estimator itself. Indeed, Bai and Fu [2] established its Bahadur efficiency.
@@
In this paper, we deal with not only the Cauchy distribution but also the Pearson Type VII distribution and our focus is the maximum likelihood estimator. Some references on the maximum likelihood estimator of the Pearson Type VII distribution are Borwein and Gabor [5], Tiku and Suresh [15], and Vaughan [16]. We provide mathematically rigorous proofs of strong consistency, asymptotic efficiency, and Bahadur efficiency for the maximum likelihood estimator. Our approach does not analyze the likelihood equation directly. We show that the asymptotic properties of the maximum likelihood estimator mirror those for the arithmetic mean of independent and identically distributed (i.i.d.) random variables with finite variance. Asymptotically, the maximum likelihood estimator for the Pearson Type VII distribution behaves well.
@@
Now we state the framework and the main result. Let $m>1/2$, which covers the heavy-tailed regime of primary interest. Let $\textup{PVII}_{m}(\mu,\sigma)$ be the Pearson Type VII distribution with location $\mu\in\mathbb{R}$ and scale $\sigma>0$. Then, the probability density function of $\textup{PVII}_{m}(\mu,\sigma)$ is given by
@@
where $c_{m}$ is the normalizing constant, specifically, $c_{m}:=\left(\int_{\mathbb{R}}(1+x^{2})^{-m}dx\right)^{-1}$. The case that $m=1$ is the Cauchy distribution.
We consider the maximum likelihood estimator of the location parameter of the Pearson Type VII distribution with known scale. We can assume that $\sigma=1$. Let $(X_{n})_{n\geq 1}$ be i.i.d. random variables on a probability space $(\Omega,\mathcal{F},P)$ following $\textup{PVII}_{m}(\theta,1)$. Let $\hat{\theta}_{n}$ be the maximum likelihood estimator of the location parameter from a sample $(X_{1},\dots,X_{n})$ of size $n$. Let $\hat{\theta}_{n}(x_{1},\dots,x_{n})$ be a measurable function on $\mathbb{R}^{n}$ which maximizes the function $\theta\mapsto\prod_{i=1}^{n}f(x_{i}-\theta)$. Such a function exists by virtue of the measurable selection theorem. Then, let $\hat{\theta}_{n}:=\hat{\theta}_{n}(X_{1},\dots,X_{n})$.
@@
$\lim_{n\to\infty}\hat{\theta}_{n}=\theta,\textup{ $P$-a.s.}$
@@
Once we see the strong consistency, it is natural to consider the asymptotic normality. We denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$ by $N(\mu,\sigma^{2})$.
@@
$\left(\sqrt{n}(\hat{\theta}_{n}-\theta)\right)_{n}$ converges to $N\left(0,\dfrac{m+1}{m(2m-1)}\right)$ in distribution as $n\to\infty$.
@@
We can also see the following behavior, although it is not often considered in mathematical statistics.
@@
$\limsup_{n\to\infty}\sqrt{\frac{n}{\log\log n}}(\hat{\theta}_{n}-\theta)=\sqrt{\frac{2(m+1)}{m(2m-1)}},\textup{ $P$-a.s.}$
@@
$\limsup_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\limsup_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\leq-\frac{m(2m-1)}{2(m+1)}.$ (1.1)
$\liminf_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\liminf_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\geq-\frac{m(2m-1)}{2(m+1)}.$ (1.2)
(ii) For every sequence $(\lambda_{n})_{n}$ of positive numbers satisfying $\lim_{n\to\infty}\lambda_{n}=\infty$ and $\lim_{n\to\infty}\lambda_{n}/n^{1/2}=0$ and every $\epsilon>0$,
$\lim_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon/\lambda_{n}\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
This assertion implies Theorem 1.1 and its proof does not depend on Theorem 1.1. However, we can show Theorem 1.1 much more easily than the proof of this assertion. For the proof, we follow the strategy of [2].
It is worth investigating the probability that the estimator deviates significantly from the true value. Hereafter, $\mathbb{N}=\{1,2,\dots\}$.
@@
There exist positive constants $c_{m},r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
$P\left(\left|\hat{\theta}_{n}-\theta\right|>r\right)\leq r^{-c_{m}n}.$
In particular, $\hat{\theta}_{n}\in L^{c_{m}n-1}(\Omega,\mathcal{F},P)$ for $n\geq N_{m}$.
@@
This is consistent with [9, (28.61c)]. We give a mathematically rigorous proof of it. The proof is technically involved and we use Theorems 1.2 and 1.5.
In the following sections, we present proofs of these assertions. In the final section, we give numerical computations of $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$.
In the proofs of these results, we can assume that $\theta=0$ without loss of generality. The parameter $m$ remains fixed throughout. Many constants will appear. When a constant depends only on $m$, we indicate this by attaching $m$ to its index; otherwise we omit it even if it depends on $m$.
@@
We prove Theorem 1.1 by following the strategy of [4, Section 3.2]. One of our goals is to establish [4, Theorem 3.3] in the case where the loss function is replaced [^1] In [4, Theorem 3.3], the loss function is given by the map $u\mapsto u^{\alpha}$ for $\alpha\in(0,1)$. with the map $u\mapsto\log(1+u^{2})$.
@@
$L_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}\log(1+(X_{i}-t)^{2}),\ t\in\mathbb{R}.$
Let $\nu_{m}$ be the Borel probability measure of the Pearson Type VII distribution $\textup{PVII}_{m}(0,1)$, that is,
$\nu_{m}(dx)=c_{m}\left(1+x^{2}\right)^{-m}dx.$
@@
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{\bf 1}_{[-1,1]}(X_{i}(\omega))=\nu_{m}([-1,1])>0,\ \textup{ $P$-a.s.$\omega$}.$
We have the assertion for $c_{m,1}:=\nu_{m}([-1,1])$. ∎
@@
$Q_{n}(\omega)=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}(w)}.$
@@
By the strong law of large numbers, there exists an event $\Omega_{2}\subset\Omega_{1}$ such that $P(\Omega_{2})=1$ and for every $\omega\in\Omega_{2}$,
@@
For a Borel probability measure $\nu$ on $\mathbb{R}$, let the expected loss function be
$F_{\nu}(t):=\int_{\mathbb{R}}\log(1+(x-t)^{2})d\nu(x),\ t\in\mathbb{R},$
and the mean set be
$C_{\nu}:=\left\{t\in\mathbb{R}\middle|\min_{s\in\mathbb{R}}F_{\nu}(s)=F_{\nu}(t)\right\}.$
For $\nu=Q_{n}(\omega)$, $C_{Q_{n}(\omega)}$ is called the Fréchet mean set. It holds that $\hat{\theta}_{n}(\omega)\in C_{Q_{n}(\omega)}$.
Another goal of this section is to show that the mean set $C_{\nu_{m}}$ is a singleton, which will be established in Lemma 2.5 below.
For notational convenience,
$F_{m}(t):=F_{\nu_{m}}(t)=\int_{\mathbb{R}}\log(1+(x-t)^{2})\nu_{m}(dx).$ (2.1)
This is the expected loss function.
@@
$P$-a.s., it holds that every $t\in\mathbb{R}$,
$\lim_{n\to\infty}L_{n}(t)=F_{m}(t).$ (2.2)
@@
By the strong law of large numbers, for every fixed $t\in\mathbb{R}$, (2.2) holds a.s. Now, use the rational approximation and use the Lipschitz continuity [^2] This estimate works well if $x$ and $y$ are close. If $x$ or $y$ is large, the bound can be very loose. of $\log(1+x^{2})$, specifically,
$\left|\log(1+x^{2})-\log(1+y^{2})\right|\leq\left||x|-|y|\right|\leq|x-y|.$ (2.3)
@@
$F_{m}$ is strictly decreasing on $(-\infty,0)$ and strictly increasing on $(0,\infty)$. In particular, $C_{\nu_{m}}=\{0\}$.
@@
$F_{m}^{\prime}(t)=-2c_{m}\int_{\mathbb{R}}\frac{x-t}{(1+(x-t)^{2})(1+x^{2})^{m}}dx.$ (2.4)
@@
Let $\varphi$ be a continuous function on $\mathbb{R}$ such that $\lim_{|z|\to\infty}\varphi(z)=\infty$. Let $C_{\varphi}:=\left\{x\in\mathbb{R}:\min_{t\in\mathbb{R}}\varphi(t)=\varphi(x)\right\}$. Then, for every $\epsilon>0$, there exists $\delta>0$ such that for every $x\in\mathbb{R}$ with $\varphi(x)\leq\min_{t\in\mathbb{R}}\varphi(t)+\delta$, $d(x,C_{\varphi})<\epsilon$.
@@
We show this by contradiction. Assume that there exists $\epsilon_{0}>0$ such that for every $n\in\mathbb{N}$, there exists $x_{n}\in\mathbb{R}$ such that $\varphi(x_{n})\leq\min_{t\in\mathbb{R}}\varphi(t)+1/n$ and $d(x_{n},C_{\varphi})\geq\epsilon_{0}$. Since $\sup_{n\in\mathbb{N}}\varphi(x_{n})<+\infty$, by the assumption of $\varphi$, $(x_{n})_{n}$ is a bounded sequence. Then, there exist a subsequence $(x_{n_{k}})_{k}$ and a point $z\in\mathbb{R}$ such that $x_{n_{k}}\to z,k\to\infty$. By the continuity of $\varphi$, $\varphi(z)=\lim_{k\to\infty}\varphi(x_{n_{k}})=\min_{t\in\mathbb{R}}\varphi(t)$. Hence, $z\in C_{\varphi}$. Now it suffices to recall that $|x_{n_{k}}-z|\geq d(x_{n_{k}},C_{\varphi})\geq\epsilon_{0}$ for each $k$. ∎
@@
Let $\epsilon>0$. Then, there exists $N_{4}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n>N_{4}(\omega,\epsilon)$ and $t_{n}\in C_{Q_{n}(\omega)}$,
@@
Now apply Lemma 2.6 to $\varphi=F_{m}$ and Lemma 2.5. ∎
By Proposition 2.7, we obtain Theorem 1.1.
@@
Recently, Schötz [13] gave precise analysis for the Fréchet mean. His approach uses a general ergodic theorem and differs from the above approach.
@@
We follow the strategy of [3, Section 4]. Let
$D(x,t):=\frac{x-t}{1+(x-t)^{2}},\ x,t\in\mathbb{R}.$
@@
$D_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}D(X_{i},t).$
Then, $2D_{n}(t)\equiv S_{n}^{\prime}(t)$ and hence, the likelihood equation is $D_{n}(t)=0$.
We first show that $\hat{\theta}_{n}>t$ holds if and only if $D_{n}(t)>0$ on an event with high probability. Since $D(x,t)$ is not monotone with respect to $t$ on $\mathbb{R}$, we cannot apply the result of [3] directly and need to “localize” it. The arguments in Section 2 are not sufficient for this goal, since the possibility that $|C_{Q_{n}}\cap[-T,T]|\geq 2$ has not yet been excluded.
Let
$G_{m}(t):=E\left[D(X_{1},t)\right]=\int_{\mathbb{R}}D(x,t)\nu_{m}(dx).$
Then, by (2.4), $2G_{m}(t)\equiv F_{m}^{\prime}(t)$ and hence, $G_{m}(-t)>0>G_{m}(t)$ for every $t>0$.
We show that
@@
$\lim_{n\to\infty}\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|=0,\ \textup{ $P$-a.s.}$
@@
Let $t\in[-1,1]$. Let $Y_{i}:=D(X_{i},t)-G_{m}(t)$. Since $|D(X_{i},t)|\leq 1/2$ and $|G_{m}(t)|\leq 1/2$, $(Y_{i})_{i}$ are i.i.d., $|Y_{i}|\leq 1$ and $E[Y_{i}]=0$. By the Azuma-Hoeffding inequality (see Petrov [11, 2.6.2]),
@@
Let $\mathcal{D}_{N}:=\{\ell/N:-N\leq\ell\leq N\}$ for $N\in\mathbb{N}$. Since $t\mapsto D(x,t)$ is Lipschitz continuous with the Lipschitz constant $1$,

### v2

@@ Abstract.
Pearson Type VII distribution, Cauchy distribution, maximum likelihood estimator, strong consistency, asymptotic normality, asymptotic efficiency, Bahadur efficiency
@@ 1. Introduction
The family of Pearson Type VII distributions provides flexible heavy-tailed models. The estimation of its parameters dates back at least to Fisher [10], over a century ago, and many researchers have studied it since then; see Johnson, Kotz, and Balakrishnan [14, Section 28] for a thorough survey of results prior to 1994. This class is also known as the location–scale family of Student’s $t$ distributions or of $q$-Gaussian distributions. For estimating the location, the median is a robust alternative to the arithmetic mean; however it is not asymptotically efficient.
@@
In general, the maximum likelihood estimator is widely regarded as optimal in large samples under standard regularity. Lange, Little, and Taylor [15] proposed a strategy based on maximum likelihood for a general model with errors following the $t$-distribution and applied it to many problems. Under suitable regularity conditions, properties such as strong consistency, asymptotic normality, and Bahadur efficiency have been established by many researchers. For location–scale families, it is natural to consider the estimation of the location with known scale. The standard approach is to solve the likelihood equation explicitly or numerically, which often has a unique root. For the Cauchy distribution with known scale, however, the likelihood equation may have multiple roots (see Reeds [17] for precise analysis), and the same phenomenon occurs for the Pearson Type VII distribution. For this reason, alternative estimators of the Cauchy location parameter have been considered. For example, Freue [11] considered the Pitman estimator for small samples, and Zhang [24] considered an empirical Bayes estimator. Nevertheless, this does not represent a failure of the maximum likelihood estimator itself. Indeed, Bai and Fu [3] established its Bahadur efficiency.
@@
In this paper, we deal with not only the Cauchy distribution but also the Pearson Type VII distribution and our focus is the maximum likelihood estimator. Some references on the maximum likelihood estimator of the Pearson Type VII distribution are Borwein and Gabor [7], Tiku and Suresh [21], and Vaughan [23]. We provide mathematically rigorous proofs of strong consistency, asymptotic efficiency, and Bahadur efficiency for the maximum likelihood estimator. Our approach does not analyze the likelihood equation directly. We show that the asymptotic properties of the maximum likelihood estimator mirror those for the arithmetic mean of independent and identically distributed (i.i.d.) random variables with finite variance. Asymptotically, the maximum likelihood estimator for the Pearson Type VII distribution performs well.
@@
Now we state the framework and the main result. Let $m>1/2$, which covers the heavy-tailed regime of primary interest. Let $\textup{PVII}_{m}(\mu,\sigma)$ be the Pearson Type VII distribution with location $\mu\in\mathbb{R}$ and scale $\sigma>0$. Then the probability density function of $\textup{PVII}_{m}(\mu,\sigma)$ is given by
@@
where $c_{m}$ is the normalizing constant, specifically, $\displaystyle c_{m}\coloneqq\left(\int_{\mathbb{R}}(1+x^{2})^{-m}dx\right)^{-1}$. The case $m=1$ corresponds to the Cauchy distribution.
We consider the maximum likelihood estimator of the location parameter of the Pearson Type VII distribution with known scale. We can assume that $\sigma=1$. Let $(X_{n})_{n\geq 1}$ be i.i.d. random variables on a complete probability space $(\Omega,\mathcal{F},P)$ following $\textup{PVII}_{m}(\theta,1)$. Let $\hat{\theta}_{n}$ be the maximum likelihood estimator of the location parameter from a sample $(X_{1},\dots,X_{n})$ of size $n$. Let $\hat{\theta}_{n}(x_{1},\dots,x_{n})$ be a measurable function on $\mathbb{R}^{n}$ which maximizes the function $\displaystyle\theta\mapsto\prod_{i=1}^{n}f(x_{i}-\theta)$. Such a function exists by virtue of the measurable selection theorem. Let $\hat{\theta}_{n}\coloneqq\hat{\theta}_{n}(X_{1},\dots,X_{n})$.
@@
$\lim_{n\to\infty}\hat{\theta}_{n}=\theta,\textup{ $\displaystyle P$-a.s.}$
@@
Once the strong consistency is given, it is natural to consider the asymptotic normality. We denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$ by $N(\mu,\sigma^{2})$.
@@
$\left(\sqrt{n}(\hat{\theta}_{n}-\theta)\right)_{n}$ converges to
$N\left(0,\dfrac{m+1}{m(2m-1)}\right)$ in distribution as $n\to\infty$.
@@
We proceed to the law of the iterated logarithm. It has connections with statistics, in particular with sequential testing. See [18, 6, 13].
@@
$\limsup_{n\to\infty}\sqrt{\frac{n}{\log\log n}}(\hat{\theta}_{n}-\theta)=\sqrt{\frac{2(m+1)}{m(2m-1)}},\textup{ $\displaystyle P$-a.s.}$
@@
(1.1) $\limsup_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\limsup_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\leq-\frac{m(2m-1)}{2(m+1)}.$
(1.2) $\liminf_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\liminf_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\geq-\frac{m(2m-1)}{2(m+1)}.$
(ii) For every sequence $(a_{n})_{n}$ of positive numbers satisfying $\displaystyle\lim_{n\to\infty}a_{n}=\infty$ and $\displaystyle\lim_{n\to\infty}a_{n}/n^{1/2}=0$ and every $\epsilon>0$,
$\lim_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon/a_{n}\right)}{n/a_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
This assertion implies Theorem 1.1 and its proof does not depend on Theorem $1.$ However we can show Theorem 1.1 much more easily than the proof of this assertion. For the proof, we follow the strategy of [3].
It is worth investigating the probability that the estimator deviates significantly from the true value. In this paper, we let $\mathbb{N}\coloneqq\{1,2,\dots\}$.
@@
There exist positive constants $r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
$P\left(\left|\hat{\theta}_{n}-\theta\right|>r\right)\leq r^{-c_{m}^{\prime}n},$
where $c_{m}^{\prime}\coloneqq\lambda_{m}^{\prime}-\dfrac{(\lambda_{m}^{\prime})^{2}}{4}$ and $\lambda_{m}^{\prime}\coloneqq\min\left\{1,\dfrac{2m-1}{4}\right\}$. In particular, $\hat{\theta}_{n}\in L^{c_{m}^{\prime}n-1}(\Omega,\mathcal{F},P)$ for $n\geq N_{m}$.
@@
This is consistent with [14, (28.61c)]. We give a mathematically rigorous proof of it. The proof is technically involved and we use Theorem 1.2 and Theorem 1.5, and an estimate obtained in the proof of Theorem 1.4.
In the following sections, we present proofs of these assertions. In the final section, we give simulation studies.
In the proofs of these results, we can assume that $\theta=0$ without loss of generality. The parameter $m$ remains fixed throughout. Many constants will appear. When a constant depends only on $m$, we indicate this by including $m$ as a subscript; otherwise we omit it even if it depends on $m$.
@@
We first give an outline of the proof. We prove Theorem 1.1 by following the strategy of Bhattacharya and Bhattacharya [5, Section 3.2]. One of our goals is to establish [5, Theorem 3.3] in the case where the loss function is replaced [^1] In [5, Theorem 3.3], the loss function is given by the map $u\mapsto u^{\alpha}$ for $\alpha\in(0,1)$. with the map $u\mapsto\log(1+u^{2})$.
The key step is Proposition 2.7, which shows that the Fréchet mean set $C_{Q_{n}(\omega)}$ (equivalently, the argmin set of $L_{n}$) is eventually contained in an arbitrarily small neighborhood of $0$. Proposition 2.7 immediately yields Theorem $1.$ The proof of Proposition 2.7 consists of four ingredients: uniform boundedness of minimizers (Lemma 2.2), a uniform law of large numbers for $L_{n}$ on bounded intervals (Lemma 2.4), uniqueness of the population minimizer $C_{\nu_{m}}=\{0\}$ (Lemma 2.5), and a stability lemma for minimizers of continuous functions with compact level sets (Lemma 2.6).
@@
$L_{n}(t)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\log(1+(X_{i}-t)^{2}),\ t\in\mathbb{R}.$
Let $\nu_{m}$ be the probability measure on the Lebesgue measurable space $(\mathbb{R},\mathcal{L}(\mathbb{R}))$ of the Pearson Type VII distribution $\textup{PVII}_{m}(0,1)$, that is,
$\nu_{m}(dx)\coloneqq c_{m}\left(1+x^{2}\right)^{-m}dx.$
@@
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{\bf 1}_{[-1,1]}(X_{i}(\omega))=\nu_{m}([-1,1])>0,\ \textup{ $\displaystyle P$-a.s. $\displaystyle\omega$}.$
We have the assertion for $c_{m,1}\coloneqq\nu_{m}([-1,1])$. ∎
@@
$Q_{n}(\omega)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}(w)}.$
For a probability measure $\nu$ on the Lebesgue measurable space $(\mathbb{R},\mathcal{L}(\mathbb{R}))$, let the expected loss function be
$F_{\nu}(t)\coloneqq\int_{\mathbb{R}}\log(1+(x-t)^{2})d\nu(x),\ t\in\mathbb{R},$
and the mean set be
$C_{\nu}\coloneqq\left\{t\in\mathbb{R}\middle|\min_{s\in\mathbb{R}}F_{\nu}(s)=F_{\nu}(t)\right\}.$
Since $\log(1+(x-t)^{2})=O\left(|x|^{\alpha}\right),x\to\infty$, for every $\alpha>0$, $F_{\nu_{m}}(t)<\infty$ for every $t\in\mathbb{R}$.
For $\nu=Q_{n}(\omega)$, $C_{Q_{n}(\omega)}$ is called the Fréchet mean set. Recall that maximizing the likelihood is equivalent to minimizing the empirical negative log-likelihood $L_{n}$. For the empirical measure $Q_{n}$, the corresponding Fréchet function $F_{Q_{n}(\omega)}(t)$ equals $L_{n}(t)(\omega)$. Therefore, $\hat{\theta}_{n}(\omega)\in C_{Q_{n}(\omega)}$.
Another goal of this section is to show that the mean set $C_{\nu_{m}}$ is a singleton, which will be established in Lemma 2.5 below. Let
(2.1) $F_{m}(t)\coloneqq F_{\nu_{m}}(t)=\int_{\mathbb{R}}\log\left(1+(x-t)^{2}\right)\nu_{m}(dx).$
This is the expected loss function.
@@
Recall (2.1). We remark that $F_{\nu_{m}}(0)<\infty$. By the strong law of large numbers, there exists an event $\Omega_{2}\subset\Omega_{1}$ such that $P(\Omega_{2})=1$ and for every $\omega\in\Omega_{2}$,
@@
$P$-a.s., it holds that for every $t\in\mathbb{R}$,
(2.2) $\lim_{n\to\infty}L_{n}(t)=F_{m}(t).$
@@
By the strong law of large numbers, for every fixed $t\in\mathbb{R}$, equation (2.2) holds a.s. More specifically, for every $t\in\mathbb{R}$, there exists an event $\Omega_{t}$ such that $P(\Omega_{t})=1$ and for every $\omega\in\Omega_{t}$, $\lim_{n\to\infty}L_{n}(t)(\omega)=F_{m}(t)$.
We use the Lipschitz continuity [^2] This estimate works well if $x$ and $y$ are close. If $x$ or $y$ is large, the bound can be very loose. of $\log(1+x^{2})$, specifically,
(2.3) $\left|\log(1+x^{2})-\log(1+y^{2})\right|\leq\left||x|-|y|\right|\leq|x-y|.$
Hence
(2.4) $\left|L_{n}(t)(\omega)-L_{n}(s)(\omega)\right|\leq|t-s|,\ \ t,s\in\mathbb{R},\,n\geq 1,$
and
(2.5) $\left|F_{m}(t)-F_{m}(s)\right|\leq|t-s|,\ t,s\in\mathbb{R}.$
We use the rational approximation. Let $\displaystyle\Omega_{\mathbb{Q}}\coloneqq\bigcap_{t\in\mathbb{Q}}\Omega_{t}$. Then $P(\Omega_{\mathbb{Q}})=1$. Take $t\in\mathbb{R}$ arbitrarily. By (2.4) and (2.5), it holds that for every $\omega\in\Omega_{\mathbb{Q}}$ and $s\in\mathbb{Q}$,
$\limsup_{n\to\infty}|L_{n}(t)(\omega)-F_{m}(t)|\leq 2|t-s|+\limsup_{n\to\infty}|L_{n}(s)(\omega)-F_{m}(s)|=2|t-s|.$
Since $s\in\mathbb{Q}$ can be taken arbitrarily close to $t$, we see that $\lim_{n\to\infty}|L_{n}(t)(\omega)-F_{m}(t)|=0$. ∎
@@
The function $F_{m}$ is strictly decreasing on $(-\infty,0)$ and strictly increasing on $(0,\infty)$. In particular, $C_{\nu_{m}}=\{0\}$.
@@
(2.6) $F_{m}^{\prime}(t)=-2c_{m}\int_{\mathbb{R}}\frac{x-t}{(1+(x-t)^{2})(1+x^{2})^{m}}dx.$
@@
Let $\varphi$ be a continuous function on $\mathbb{R}$ such that $\displaystyle\lim_{|z|\to\infty}\varphi(z)=\infty$. Let $\displaystyle C_{\varphi}\coloneqq\left\{x\in\mathbb{R}:\min_{t\in\mathbb{R}}\varphi(t)=\varphi(x)\right\}$. Then, for every $\epsilon>0$, there exists $\delta>0$ such that for every $x\in\mathbb{R}$ with $\displaystyle\varphi(x)\leq\min_{t\in\mathbb{R}}\varphi(t)+\delta$, $d(x,C_{\varphi})<\epsilon$.
@@
We show this by contradiction. Assume that there exists $\epsilon_{0}>0$ such that for every $n\in\mathbb{N}$, there exists $x_{n}\in\mathbb{R}$ such that $\varphi(x_{n})\leq\min_{t\in\mathbb{R}}\varphi(t)+1/n$ and $d(x_{n},C_{\varphi})\geq\epsilon_{0}$. Since $\displaystyle\sup_{n\in\mathbb{N}}\varphi(x_{n})<+\infty$, by the assumption of $\varphi$, $(x_{n})_{n}$ is a bounded sequence. Then there exist a subsequence $(x_{n_{k}})_{k}$ and a point $z\in\mathbb{R}$ such that $x_{n_{k}}\to z,k\to\infty$. By the continuity of $\varphi$, $\displaystyle\varphi(z)=\lim_{k\to\infty}\varphi(x_{n_{k}})=\min_{t\in\mathbb{R}}\varphi(t)$. Hence, $z\in C_{\varphi}$. Now it suffices to recall that $|x_{n_{k}}-z|\geq d(x_{n_{k}},C_{\varphi})\geq\epsilon_{0}$ for each $k$. ∎
@@
Let $\epsilon>0$. Then there exists $N_{4}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n>N_{4}(\omega,\epsilon)$ and $t_{n}\in C_{Q_{n}(\omega)}$,
@@
Now apply Lemma 2.6 to $\displaystyle\varphi=F_{m}$ and Lemma 2.5. ∎
Recall that $\hat{\theta}_{n}$ is a measurable selection from $C_{Q_{n}(\omega)}$. By Proposition 2.7, we obtain Theorem $1.$
@@
Recently, Schötz [19] gave a precise analysis for the Fréchet mean. His approach uses a general ergodic theorem and differs from the above approach.
@@
We first give an outline of the proof. We follow the strategy of Barczy and Páles [4, Section 4]. Let
$D(x,t)\coloneqq\frac{x-t}{1+(x-t)^{2}},\ x,t\in\mathbb{R},$
and
$D_{n}(t)\coloneqq\frac{1}{n}\sum_{i=1}^{n}D(X_{i},t),\ t\in\mathbb{R}.$
Then $-2D_{n}(t)\equiv L_{n}^{\prime}(t)$ and hence, the likelihood equation is $D_{n}(t)=0$.
Since the map $t\mapsto D(x,t)$ is not monotone on $\mathbb{R}$ for each fixed $x$, we cannot apply the result of [4, Section 4] directly. Therefore we “localize” the argument. Specifically, we construct a sequence of events $\mathcal{A}_{n},n\geq 1$, defined in (3.2) below, such that $\displaystyle\lim_{n\to\infty}P(\mathcal{A}_{n})=1$, and, on each $\mathcal{A}_{n}$ the map $t\mapsto D_{n}(t)$ is strictly decreasing on a fixed interval $I$ and has a unique zero in $I$. This yields that on each $\mathcal{A}_{n}$, $\hat{\theta}_{n}>t$ if and only if $D_{n}(t)>0$ for $t\in I$. Once this localization is established, the remainder of the proof is standard. Consider a Taylor expansion of $D_{n}(y/\sqrt{n})$ at $0$ and apply the central limit theorem, the law of large numbers and Slutsky’s lemma.
The arguments in Section 2 are not sufficient for this localization, since the possibility that $|C_{Q_{n}}\cap[-T,T]|\geq 2$ has not yet been excluded. We overcome this issue by Proposition 3.4. For the proof, we compare $D_{n}(t)$ and its derivative $D_{n}^{\prime}(t)$ with their population counterparts $G_{m}(t)$ and $G_{m}^{\prime}(t)$ defined below. Lemma 3.1 shows that $D_{n}(t)$ is uniformly close to $G_{m}(t)$ for large $n$. Lemma 3.2 shows that $G_{m}^{\prime}(t)<0$. Lemma 3.3 transfers this to $D_{n}^{\prime}(t)$ for large $n$.
Theorem 1.2 can also be shown by using van der Vaart [22, Theorem 5.23]. See Remark 3.5 (iii) below for details. However, throughout this paper we repeatedly use the notation such as $\{\mathcal{A}_{n}\}_{n\geq 1}$ and refer to related assertions in this section, we therefore present the full details here.
@@
$G_{m}(t)\coloneqq E\left[D(X_{1},t)\right]=\int_{\mathbb{R}}D(x,t)\nu_{m}(dx).$
Then, by (2.6), $-2G_{m}(t)\equiv F_{m}^{\prime}(t)$ and hence, $G_{m}(-t)>0>G_{m}(t)$ for every $t>0$.
We show the following:
@@
$\lim_{n\to\infty}\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|=0,\ \textup{ $\displaystyle P$-a.s.}$
@@
Let $t\in[-1,1]$. Let $Y_{i}\coloneqq D(X_{i},t)-G_{m}(t)$. Since $|D(X_{i},t)|\leq 1/2$ and $|G_{m}(t)|\leq 1/2$, it holds that $(Y_{i})_{i}$ are i.i.d., $|Y_{i}|\leq 1$ and $E[Y_{i}]=0$. By the Azuma-Hoeffding inequality (see Petrov [16, 2.6.2] or Boucheron, Lugosi and Massart [8, Theorem 2.8]),
@@
Let $\mathcal{D}_{N}\coloneqq\{\ell/N:-N\leq\ell\leq N\}$ for $N\in\mathbb{N}$. Since $t\mapsto D(x,t)$ is Lipschitz continuous with the Lipschitz constant $1$,